# LAPS-Diff: A Diffusion-Based Framework for Singing Voice Synthesis With Language Aware Prosody-Style Guided Learning

### v1

@@ Abstract
The field of Singing Voice Synthesis (SVS) has seen significant advancements in recent years due to the rapid progress of diffusion-based approaches. However, capturing vocal style, genre-specific pitch inflections, and language-dependent characteristics remains challenging, particularly in low-resource scenarios. To address this, we propose LAPS-Diff, a diffusion model integrated with language-aware embeddings and a vocal-style guided learning mechanism, specifically designed for Bollywood Hindi singing style. We curate a Hindi SVS dataset and leverage pre-trained language models to extract word and phone-level embeddings for an enriched lyrics representation. Additionally, we incorporated a style encoder and a pitch extraction model to compute style and pitch losses, capturing features essential to the naturalness and expressiveness of the synthesized singing, particularly in terms of vocal style and pitch variations. Furthermore, we utilize MERT and IndicWav2Vec models to extract musical and contextual embeddings, serving as conditional priors to refine the acoustic feature generation process further. Based on objective and subjective evaluations, we demonstrate that LAPS-Diff significantly improves the quality of the generated samples compared to the considered state-of-the-art (SOTA) model for our constrained dataset that is typical of the low resource scenario.
Index Terms:
Singing Voice Synthesis, Diffusion Model, Bollywood Hindi Singing Style.
@@ I Introduction
Singing Voice Synthesis (SVS) is the process of generating natural-sounding singing voices using statistical or Artificial Intelligence (AI)-based algorithms, guided by musical scores [1]. Unlike Text-to-Speech (TTS) synthesis systems that convert text into audible speech [2], the primary objective of SVS models is to produce a singing voice that is accurately synchronized with the musical composition as given by a music score. Like a TTS model, an SVS model also follows a three-stage pipeline: converting the input music score into bottleneck features, generating acoustic features from bottleneck features, and reconstructing the audible singing voice [3]. In recent years, the growing use of SVS applications in the media and entertainment industry has highlighted the significant potential of this evolving field [2, 4].
@@
In the recent past, improved variants of diffusion models have exhibited significant advancements in SVS [5, 6, 7, 8]. In particular, the DiffSinger model [1] is widely recognized as a pioneering approach in diffusion-based SVS research and has demonstrated promising results. It includes an auxiliary decoder that generates the target mel-spectrogram from music score embeddings. Specifically, the shallow diffusion mechanism introduced in DiffSinger accelerates the inference by providing the denoiser a boundary predicted intermediate mel-spectrogram rather than the noise input. Learn2Sing 2.0 [2] is also a well-known diffusion-based approach in which the authors adapted the GradTTS framework [9] to transform speech data into singing audio. In Learn2Sing 2.0, the authors effectively demonstrated how TTS models can be adapted for SVS with significant adjustments. Specifically, to improve the denoising process in the reverse diffusion for accurate estimation of the denoised mel-spectrogram having precise detail of the speaker’s vocal characteristics, Learn2Sing 2.0 [2] introduced style embedding as an additional conditioning feature within the Grad-TTS framework. As a newer variant of diffusion-based approach, the authors of MakeSinger [5] introduced an efficient semi-supervised training mechanism, enabling the model to learn from both labeled and unlabeled singing data. They conditioned the denoiser on text, pitch, and speaker embeddings, ensuring an improved reverse denoising process. While diffusion-based approaches have shown impressive results, models like VISinger2 [10] have also exhibited notable performance in SVS by utilizing a conditional variational autoencoder (CVAE)-based framework.
@@
Although the discussed models have shown significant advancements, most of these existing SVS models, including DiffSinger, have been trained and evaluated on singing audio datasets [11, 12, 13, 14] that typically contain more than 5 hours of audio data (single voice) in languages such as Mandarin, Korean, Japanese, and English. Moreover, each of these language-specific singing datasets belongs to a distinct musical genre with its own peculiar characteristics for melody (pitch variation), rhythm, and pronunciation. These aspects are also strongly influenced by the individual singer’s unique vocal characteristics. However, due to the scarcity of labeled singing data in most languages and genres, it is highly challenging for an SVS model to effectively capture linguistic content, style, and pitch related information from low-resource singing data, highlighting a significant research gap.
@@
To address this challenge of capturing content information, vocal style, and pitch variations from low-resource SVS data, we introduce LAPS-Diff, a diffusion model that incorporates language-aware embeddings and prosody-style guided learning mechanism. The proposed model is built on the DiffSinger framework [1]. As part of this work, we curate a dataset of about one hour duration of Hindi Bollywood-style songs by a single male singer and process the audio to obtain the music scores. To effectively capture the lyrical content information, we leverage additionally two pre-trained language models, IndicBERT [15] and XPhoneBERT [16], for Hindi word and phone-level embeddings. We combine these embeddings with music score embedding [1] to form an enriched content representation. Further, we integrate a style encoder [17] and pre-trained JDCNet pitch extraction model [18, 19] to obtain style and pitch information (melody) to compute the corresponding losses for the training of the auxilary decoder of our proposed LAPS-Diff model. Given the critical importance of preserving expressive pitch inflections in the songs in the genre of interest, we go beyond L1 loss of F0 (or log F0), and investigate the role of linear correlation in the loss function. The resulting style and pitch losses are found to improve the model’s ability to effectively capture vocal style and pitch-related information, enabling a closer resemblance to the underlying singing dynamics. Moreover, we utilize pre-trained MERT [20] and IndicWav2Vec [21] models for extracting musical feature embeddings and contextual embeddings, respectively, to use as conditional priors to the denoiser for improving the reverse diffusion process [22] through explicit feature guidance. Finally, we evaluate the performances of the proposed LAPS-Diff model and the considered state-of-the-art (SOTA) SVS model using our Bollywood Hindi dataset with objective and subjective measures and discuss the obtained improvements.
@@ II Dataset
The original song audio (as obtained from publicly accessible sites) is processed for vocals separation using a commercial tool [23]. Automatically detected silences greater than 500 ms are used to segment the song into the lyrics lines/phrases that are then manually assigned to each audio segment. Next, we achieve the automatic alignment of the audio and text at the phone level using forced alignment with a hybrid automatic speech recognition system [24] that has been trained on 180 hours of adult Hindi read speech across 400 male and female speakers [25]. Sung lyrics exhibit notable differences from speech, chiefly on word pronunciations in terms of vowel duration as well as the nature of pitch variation over this duration. It is also common to find the addition of schwa on the last consonant of words. We account for these differences with a specially constructed lexicon by the systematic augmentation with alternate pronunciations. Syllable boundaries are obtained by merging the corresponding aligned phones. The vocal pitch is extracted at 10 ms intervals using an autocorrelation based method for fundamental frequency and voicing [26]. Brief uv segments and pauses are linearly interpolated in pitch. Each syllable is assigned a MIDI note corresponding to the quantized mode of the F0 values across the syllable segment. Syllables longer than 200 ms are analysed for pitch fluctuations by computing the MIDI pitch separately for non-overlapping sub-segments of 200 ms and assigning a Slur flag to those segments that register a MIDI note change from the previous. This is the first attempt to the best of our knowledge to build an Indian music dataset for an SVS task. We adopt the same music score format as the Opencpop dataset [11] with lyrical content in text, corresponding phoneme sequence, associated musical notes, note duration, phoneme duration, and slur information, as depicted in Figure 1.
@@ III Proposed Method
The proposed model is built upon the DiffSinger framework [1]. The same encoder model $E(.)$ is utilized to extract the music score embedding ${\boldsymbol{e}_{m}}$ from the music score $x$, as depicted in Fig. 2. Similar to DiffSinger, the LAPS-Diff model incorporates two key components: the auxiliary decoder (aux decoder) and the denoiser. Training details of the auxiliary decoder and the denoiser, as well as the inference stage, are discussed in the following subsections.
@@ III-A Feature Fusion
To efficiently extract Hindi lyrical content related information from the textual content of the music score by obtaining an enriched lyrical representation, we use two pre-trained language models: IndicBERT (as a text encoder $T(.)$) and XPhoneBERT (as a phone encoder $H(.)$) both of which are compatible with the Hindi language. Thereafter, we combine IndicBERT’s word level embedding ${\boldsymbol{e}_{w}}$ and XPhoneBERT’s phone level embedding ${\boldsymbol{e}_{p}}$ with ${\boldsymbol{e}_{m}}$ to derive the final fused embedding ${\boldsymbol{e}_{c}}$ (i.e., fusion of three feature embeddings via summation operation [27, 28]), as described mathematically in Eq.(1) (shown in Fig. 2),
@@ III-B Decoder Training
The auxiliary decoder $AD(.)$ used in this work is a mel-spectrogram decoder capable of reconstructing mel-spectrograms $\hat{M}$ from the music score embedding [1]. However, in this work, we provide the combined embedding ${\boldsymbol{e}_{c}}$ as the input to the auxiliary decoder to generate mel-spectrograms (i.e., $\hat{M}=AD(\boldsymbol{e}_{c})$, where $\hat{M}$ is the generated version that is conditioned on the supplied embedding. In addition to content information, singer’s vocal style and pitch variations carry the recognizable characteristics associated with any genre of singing. Hence, to effectively train the auxiliary decoder in capturing relevant information, we incorporate style and pitch losses as optimization objectives in our work.
@@ III-B1 Style Loss
To extract the vocal style related information in terms of style embedding (style vector) from mel-spectrogram, we used a residual connection based Convolutional Neural Network (CNN) model as the style encoder $S(.)$ (the architectural framework of $S(.)$ is similar to the acoustic style encoder of StyleTTS2 [17]). The extracted style embedding carries an abstract representation (i.e., latent representation) of the singer’s vocal characteristics. To compute the style loss $\mathcal{L}_{{style}}$, we first extract the style embeddings $\boldsymbol{e}_{s_{m}}$ and $\boldsymbol{e}_{s_{\hat{m}}}$ from both $M$ and $\hat{M}$ respectively, as illustrated in Fig. 2. We then obtain $\mathcal{L}_{{style}}$ using the Mean Squared Error (MSE), as defined in the following equation,
@@
where $N$ is the total length of the embedding. In Eq.(2), $\boldsymbol{e}_{s_{m}}=S(M)$ and $\boldsymbol{e}_{s_{\hat{m}}}=S(\hat{M})$.
@@ III-B2 Pitch Loss
The pitch loss $\mathcal{L}_{{pitch}}$ is calculated using MSE between the pitch values ${{F0}}_{m}$ and ${{F0}}_{\hat{m}}$ extracted from both $M$ and $\hat{M}$ respectively, using the pre-trained JDCNet model [19] denoted as $J(.)$. The pre-trained JDCNet model used in this work is adopted from StyleTTS2 [17]). Due to StyleTTS2’s impressive voice cloning abilities in terms of naturalness and expressiveness, we kept the framework of $S(.)$ and $J(.)$ the same as in StyleTTS2. Moreover, JDCNet [19] is extensively trained on singing voice data, making it highly effective in capturing fine pitch details from mel-spectrograms [18]. In a novel approach in this work, we introduce the concordance correlation coefficient (CCC) [29] to exploit the linear correlation between matched pitch contours ${{F0}}_{m}$ and ${{F0}}_{\hat{m}}$. That is, the CCC (given in the Eq.(3) below) balances the requirements of pitch height and correlation, making it particularly suitable for evaluating the perceptual similarity of two pitch contours. The CCC value ranges from $-1$ to $1$. To influence the pitch loss, $\mathcal{L}_{{pitch}}$, we incorporate $(1-\text{CCC})$ as a multiplicative factor such that a highly correlated pair of ${{F0}}_{m}$ and ${F0}_{\hat{m}}$ results in a lower penalty, whereas a lower correlation increases the penalty. The pitch loss $\mathcal{L}_{{pitch}}$ is mathematically defined in Eq.(3),
@@
and $\rho$ denotes the Pearson correlation coefficient between the predicted and ground-truth pitch contours, $\sigma$ represents the standard deviation, and $\mu$ denotes the mean value of the respective sequences. In Eq.(3), $K$ represents the length of the $F0$ sequence, ${{F0}}_{m}=J(M)$ and ${{F0}}_{\hat{m}}=J(\hat{M})$. To prevent the loss from becoming zero for highly correlated pairs, we introduce a minimum multiplicative factor of $0.01$.
Besides $\mathcal{L}_{{style}}$ and $\mathcal{L}_{{pitch}}$, we incorporate $L_{1}$ loss (MAE loss) to measure the difference between $M$ and $\hat{M}$, following a similar approach to DiffSinger. Apart from these three losses, all other losses remain the same as in DiffSinger’s Github repository [30].
@@ III-C Denoiser Training
In the training phase of the denoiser $D(.)$ in DiffSinger, the model takes a noisy mel-spectrogram $M_{t}$ from the $t$-th ($t\in{0,T}$) diffusion step, along with the step embedding $\boldsymbol{e}_{t}$ for time step $t$ and the music score embedding $\boldsymbol{e}_{m}$, to estimate the noise $\hat{\epsilon}_{\theta}(\cdot)$ [1] for obtaining $(t-1)$-th denoised melspectrogram $M_{t-1}$ as given in Eq.(4),
@@
In Eq.(4), $\alpha_{t}=1-\beta_{t}$ ($\beta_{t}$ is variance schedule at time step $t$), $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$, $z\sim\mathcal{N}(0,I)$ and $\sigma_{t}$ is noise scale at time step $t$ [22].
To improve the denoising process in reverse diffusion, particularly for accurate estimation of the denoised mel-spectrogram having precise detail of the musical and linguistic information, we introduce two additional feature embeddings as conditional priors to provide more explicit guidance during denoising. Between the two, one captures musical characteristics through MERT embedding [20] (focusing on tonal and pitch-related attributes), and the other captures linguistic information in terms of IndicWav2Vec [21] embedding to obtain a richer representation in the denoised mel-spectrogram. MERT and IndicWav2Vec are two pre-trained models (trained using self-supervised learning approach) that efficiently extract respective embeddings from singing audio data (shown in Fig. 2), as they are trained on singing (music) data and multilingual speech data (including Hindi), respectively. Therefore, Eq.(4) can be reformulated as follows:
@@
In Eq.(III-C), $\boldsymbol{e}_{{mert}}$ and $\boldsymbol{e}_{{vec}}$ represents MERT and IndicWav2Vec embeddings respectively. The denoiser training loss used in our work is identical to that of DiffSinger [30] (i.e., $L_{1}$ loss). The model architecture of $D(.)$ considered in LAPS-Diff is identical to the denoiser used in DiffSinger [1].

### v2

@@ Abstract
The field of Singing Voice Synthesis (SVS) has seen significant advancements in recent years due to the rapid progress of diffusion-based approaches. However, capturing vocal style, genre-specific pitch inflections, and language-dependent characteristics remains challenging, particularly in low-resource scenarios. To address this, we propose LAPS-Diff, a diffusion model integrated with language-aware embeddings and a vocal-style guided learning mechanism, specifically designed for Bollywood Hindi singing style. We curate a Hindi SVS dataset and leverage pre-trained language models to extract word and phone-level embeddings for an enriched lyrics representation. Additionally, we incorporated a style encoder and a pitch extraction model to compute style and pitch losses, capturing features essential to the naturalness and expressiveness of the synthesized singing, particularly in terms of vocal style and pitch variations. Furthermore, we utilize MERT and IndicWav2Vec models to extract musical and contextual embeddings, serving as conditional priors to refine the acoustic feature generation process further. Based on objective and subjective evaluations, we demonstrate that LAPS-Diff significantly improves the quality of the generated samples compared to the considered state-of-the-art (SOTA) model for our constrained dataset that is typical of the low resource scenario.111* Sandipan Dhar and Mayank Gupta are co-first authors
@@ I Introduction
Singing Voice Synthesis (SVS) is the process of generating natural-sounding singing voices using statistical or Artificial Intelligence (AI)-based algorithms, guided by musical scores [Ref1-Diffsinger]. Unlike Text-to-Speech (TTS) synthesis systems that convert text into audible speech [Ref2-Learn2Sing], the primary objective of SVS models is to produce a singing voice that is accurately synchronized with the musical composition as given by a music score. Like a TTS model, an SVS model also follows a three-stage pipeline: converting the input music score into bottleneck features, generating acoustic features from bottleneck features, and reconstructing the audible singing voice [Ref3-TCSinger]. In recent years, the growing use of SVS applications in the media and entertainment industry has highlighted the significant potential of this evolving field [Ref2-Learn2Sing, Ref1-Singing-Tacotron-acm-2].
@@
In the recent past, improved variants of diffusion models have exhibited significant advancements in SVS [Ref4-MakeSinger, Ref5-HiddenSinger, Ref6-RDSinger, Ref1-Singing-ExpressiveSinger-acm-3]. In particular, the DiffSinger model [Ref1-Diffsinger] is widely recognized as a pioneering approach in diffusion-based SVS research and has demonstrated promising results. It includes an auxiliary decoder that generates the target mel-spectrogram from music score embeddings. Specifically, the shallow diffusion mechanism introduced in DiffSinger accelerates the inference by providing the denoiser a boundary predicted intermediate mel-spectrogram rather than the noise input. Learn2Sing 2.0 [Ref2-Learn2Sing] is also a well-known diffusion-based approach in which the authors adapted the GradTTS framework [Ref19-Grad-TTS] to transform speech data into singing audio. In Learn2Sing 2.0, the authors effectively demonstrated how TTS models can be adapted for SVS with significant adjustments. Specifically, to improve the denoising process in the reverse diffusion for accurate estimation of the denoised mel-spectrogram having precise detail of the speaker’s vocal characteristics, Learn2Sing 2.0 [Ref2-Learn2Sing] introduced style embedding as an additional conditioning feature within the Grad-TTS framework. As a newer variant of diffusion-based approach, the authors of MakeSinger [Ref4-MakeSinger] introduced an efficient semi-supervised training mechanism, enabling the model to learn from both labeled and unlabeled singing data. They conditioned the denoiser on text, pitch, and speaker embeddings, ensuring an improved reverse denoising process. While diffusion-based approaches have shown impressive results, models like VISinger2 [Ref20-VISinger2] have also exhibited notable performance in SVS by utilizing a conditional variational autoencoder (CVAE)-based framework.
@@
Although the discussed models have shown significant advancements, most of these existing SVS models, including DiffSinger, have been trained and evaluated on singing audio datasets [Ref22-Opencpop, Ref21-Korean-Dataset, Ref24-JVS-MuSiC-Japanese, Ref23-Bisinger-English] that typically contain more than 5 hours of audio data (single voice) in languages such as Mandarin, Korean, Japanese, and English. Moreover, each of these language-specific singing datasets belongs to a distinct musical genre with its own peculiar characteristics for melody (pitch variation), rhythm, and pronunciation. These aspects are also strongly influenced by the individual singer’s unique vocal characteristics. However, due to the scarcity of labeled singing data in most languages and genres, it is highly challenging for an SVS model to effectively capture linguistic content, style, and pitch related information from low-resource singing data, highlighting a significant research gap.
@@
To address this challenge of capturing content information, vocal style, and pitch variations from low-resource SVS data, we introduce LAPS-Diff, a diffusion model that incorporates language-aware embeddings and prosody-style guided learning mechanism. The proposed model is built on the DiffSinger framework [Ref1-Diffsinger]. As part of this work, we curate a dataset of about one hour duration of Hindi Bollywood-style songs by a single male singer and process the audio to obtain the music scores. To effectively capture the lyrical content information, we leverage additionally two pre-trained language models, IndicBERT [Ref11-IndicBERT] and XPhoneBERT [Ref12-XPhoneBERT], for Hindi word and phone-level embeddings. We combine these embeddings with music score embedding [Ref1-Diffsinger] to form an enriched content representation. Further, we integrate a style encoder [Ref14-StyleTTS2] and pre-trained JDCNet pitch extraction model [Ref13-StyleTTS, Ref15-JDCNet] to obtain style and pitch information (melody) to compute the corresponding losses for the training of the auxilary decoder of our proposed LAPS-Diff model. Given the critical importance of preserving expressive pitch inflections in the songs in the genre of interest, we go beyond L1 loss of F0 (or log F0), and investigate the role of linear correlation in the loss function. The resulting style and pitch losses are found to improve the model’s ability to effectively capture vocal style and pitch-related information, enabling a closer resemblance to the underlying singing dynamics. Moreover, we utilize pre-trained MERT [Ref17-MERT] and IndicWav2Vec [Ref18-Indicwav2vec] models for extracting musical feature embeddings and contextual embeddings, respectively, to use as conditional priors to the denoiser for improving the reverse diffusion process [Ref25-Diffusion-Process] through explicit feature guidance. Finally, we evaluate the performances of the proposed LAPS-Diff model and the considered state-of-the-art (SOTA) SVS model using our Bollywood Hindi dataset with objective and subjective measures and discuss the obtained improvements.
@@ II Dataset
The original song audio (as obtained from publicly accessible sites) is processed for vocals separation using a commercial tool [Ref-GaudioLab]. Automatically detected silences greater than 500 ms are used to segment the song into the lyrics lines/phrases that are then manually assigned to each audio segment. Next, we achieve the automatic alignment of the audio and text at the phone level using forced alignment with a hybrid automatic speech recognition system [povey2011kaldi] that has been trained on 180 hours of adult Hindi read speech across 400 male and female speakers [IITM_ASR_Challenge]. Sung lyrics exhibit notable differences from speech, chiefly on word pronunciations in terms of vowel duration as well as the nature of pitch variation over this duration. It is also common to find the addition of schwa on the last consonant of words. We account for these differences with a specially constructed lexicon by the systematic augmentation with alternate pronunciations. Syllable boundaries are obtained by merging the corresponding aligned phones. The vocal pitch is extracted at 10 ms intervals using an autocorrelation based method for fundamental frequency and voicing [jadoul2018introducing]. Brief uv segments and pauses are linearly interpolated in pitch. Each syllable is assigned a MIDI note corresponding to the quantized mode of the F0 values across the syllable segment. Syllables longer than 200 ms are analysed for pitch fluctuations by computing the MIDI pitch separately for non-overlapping sub-segments of 200 ms and assigning a Slur flag to those segments that register a MIDI note change from the previous. This is the first attempt to the best of our knowledge to build an Indian music dataset for an SVS task. We adopt the same music score format as the Opencpop dataset [Ref22-Opencpop] with lyrical content in text, corresponding phoneme sequence, associated musical notes, note duration, phoneme duration, and slur information, as depicted in Figure 1.
@@ III Proposed Method
The proposed model is built upon the DiffSinger framework [Ref1-Diffsinger]. The same encoder model $E(.)$ is utilized to extract the music score embedding $\boldsymbol{e}_{m}$} from the music score $x$, as depicted in Fig. 2. Similar to DiffSinger, the LAPS-Diff model incorporates two key components: the auxiliary decoder (aux decoder) and the denoiser. Training details of the auxiliary decoder and the denoiser, as well as the inference stage, are discussed in the following subsections.
@@ III-A Feature Fusion
To efficiently extract Hindi lyrical content related information from the textual content of the music score by obtaining an enriched lyrical representation, we use two pre-trained language models: IndicBERT (as a text encoder $T(.)$) and XPhoneBERT (as a phone encoder $H(.)$) both of which are compatible with the Hindi language. Thereafter, we combine IndicBERT’s word level embedding $\boldsymbol{e}_{w}$} and XPhoneBERT’s phone level embedding $\boldsymbol{e}_{p}$} with $\boldsymbol{e}_{m}$} to derive the final fused embedding $\boldsymbol{e}_{c}$} (i.e., fusion of three feature embeddings via summation operation [Ref33-Summation-of-embeddings, Ref34-Summation-of-embeddings-1]), as described mathematically in Eq.(1) (shown in Fig. 2),
@@ III-B Decoder Training
The auxiliary decoder $AD(.)$ used in this work is a mel-spectrogram decoder capable of reconstructing mel-spectrograms $\hat{M}$ from the music score embedding [Ref1-Diffsinger]. However, in this work, we provide the combined embedding $\boldsymbol{e}_{c}$} as the input to the auxiliary decoder to generate mel-spectrograms (i.e., $\hat{M}=AD(\boldsymbol{e}_{c})$, where $\hat{M}$ is the generated version that is conditioned on the supplied embedding. In addition to content information, singer’s vocal style and pitch variations carry the recognizable characteristics associated with any genre of singing. Hence, to effectively train the auxiliary decoder in capturing relevant information, we incorporate style and pitch losses as optimization objectives in our work.
@@ III-B1 Style Loss
To extract the vocal style related information in terms of style embedding (style vector) from mel-spectrogram, we used a residual connection based Convolutional Neural Network (CNN) model as the style encoder $S(.)$ (the architectural framework of $S(.)$ is similar to the acoustic style encoder of StyleTTS2 [Ref14-StyleTTS2]). The extracted style embedding carries an abstract representation (i.e., latent representation) of the singer’s vocal characteristics. To compute the style loss $\mathcal{L}_{{style}}$, we first extract the style embeddings $\boldsymbol{e}_{s_{m}}$ and $\boldsymbol{e}_{s_{\hat{m}}}$ from both $M$ and $\hat{M}$ respectively, as illustrated in Fig. 2. We then obtain $\mathcal{L}_{{style}}$ using the Mean Squared Error (MSE), as defined in the following equation,
@@
[…]
@@ III-B2 Pitch Loss
The pitch loss $\mathcal{L}_{{pitch}}$ is calculated using MSE between the pitch values ${{F0}}_{m}$ and $^{{F0}}_{\hat{m}}$ extracted from both $M$ and $\hat{M}$ respectively, using the pre-trained JDCNet model [Ref15-JDCNet] denoted as $J(.)$. The pre-trained JDCNet model used in this work is adopted from StyleTTS2 [Ref14-StyleTTS2]). Due to StyleTTS2’s impressive voice cloning abilities in terms of naturalness and expressiveness, we kept the framework of $S(.)$ and $J(.)$ the same as in StyleTTS2. Moreover, JDCNet [Ref15-JDCNet] is extensively trained on singing voice data, making it highly effective in capturing fine pitch details from mel-spectrograms [Ref13-StyleTTS]. In a novel approach in this work, we introduce the concordance correlation coefficient (CCC) [Ref34-correlation-coefficient] to exploit the linear correlation between matched pitch contours ${{F0}}_{m}$ and $^{{F0}}_{\hat{m}}$. That is, the CCC (given in the Eq.(3) below) balances the requirements of pitch height and correlation, making it particularly suitable for evaluating the perceptual similarity of two pitch contours. The CCC value ranges from 1-1 to $1$. To influence the pitch loss, $\mathcal{L}_{{pitch}}$, we incorporate $(1-\text{CCC})$ as a multiplicative factor such that a highly correlated pair of ${{F0}}_{m}$ and $^{F0}_{\hat{m}}$ results in a lower penalty, whereas a lower correlation increases the penalty. The pitch loss $\mathcal{L}_{{pitch}}$ is mathematically defined in Eq.(3),
@@
[…]
Besides $\mathcal{L}_{{style}}$ and $\mathcal{L}_{{pitch}}$, we incorporate $L_{1}$ loss (MAE loss) to measure the difference between $M$ and $\hat{M}$, following a similar approach to DiffSinger. Apart from these three losses, all other losses remain the same as in DiffSinger’s Github repository [DiffSinger-code].
@@ III-C Denoiser Training
In the training phase of the denoiser $D(.)$ in DiffSinger, the model takes a noisy mel-spectrogram $M_{t}$ from the $t$-th ($t\in{0,T}$) diffusion step, along with the step embedding $\boldsymbol{e}_{t}$ for time step $t$ and the music score embedding $\boldsymbol{e}_{m}$, to estimate the noise $\hat{\epsilon}_{\theta}(\cdot)$ [Ref1-Diffsinger] for obtaining (t1)(t-1)-th denoised melspectrogram $M_{t-1}$ as given in Eq.(4),
@@
In Eq.(4), $\alpha_{t}=1-\beta_{t}$ ($\beta_{t}$ is variance schedule at time step $t$), $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$, $z\sim\mathcal{N}(0,I)$ and $\sigma_{t}$ is noise scale at time step $t$ [Ref25-Diffusion-Process].
To improve the denoising process in reverse diffusion, particularly for accurate estimation of the denoised mel-spectrogram having precise detail of the musical and linguistic information, we introduce two additional feature embeddings as conditional priors to provide more explicit guidance during denoising. Between the two, one captures musical characteristics through MERT embedding [Ref17-MERT] (focusing on tonal and pitch-related attributes), and the other captures linguistic information in terms of IndicWav2Vec [Ref18-Indicwav2vec] embedding to obtain a richer representation in the denoised mel-spectrogram. MERT and IndicWav2Vec are two pre-trained models (trained using self-supervised learning approach) that efficiently extract respective embeddings from singing audio data (shown in Fig. 2), as they are trained on singing (music) data and multilingual speech data (including Hindi), respectively. Therefore, Eq.(4) can be reformulated as follows:
@@
In Eq.(III-C), $\boldsymbol{e}_{{mert}}$ and $\boldsymbol{e}_{{vec}}$ represents MERT and IndicWav2Vec embeddings respectively. The denoiser training loss used in our work is identical to that of DiffSinger [DiffSinger-code] (i.e., $L_{1}$ loss). The model architecture of $D(.)$ considered in LAPS-Diff is identical to the denoiser used in DiffSinger [Ref1-Diffsinger].